# The sup-norm problem for newforms of large level on $\operatorname{PGL}(n)$

### v1

@@ 6.1. An overview
$z=\begin{pmatrix}\sqrt{y}&x/\sqrt{y}\\ 0&1/\sqrt{y}\end{pmatrix}$
and let $\gamma\in H(z,m,N)$. The bound
$z^{-1}\gamma z=O(m^{1/2})$ (6.1)
@@
for $i=1,2$, where $B(r)$ is a Euclidean ball of radius $O(r)$ around 00.
We assume now that $z$ lies in what we call the balanced bulk, as in the second alternative in Proposition 5.2, meaning that $z$ reduces to some compact $\Omega$ and $y\asymp 1/N$. Let
@@
which defines a sublattice of index $N$ of $L_{z}$. We prove in Lemma 6.1 that the lattices defined by $z$, $z^{\prime}$, and $z_{N}$ are balanced. That is to say that their respective successive minima and covolume satisfy
$\lambda_{1}\asymp\lambda_{2}\asymp\sqrt{\operatorname{vol}}.$
Notice also that
$\left\lVert e_{2}\right\rVert_{z}=\left\lVert(0,1/\sqrt{y})\right\rVert\asymp\sqrt{N}\asymp\sqrt{\operatorname{vol}(L_{z_{N}})},$
making $e_{2}\cdot z$ one of the shortest vectors in $L_{z_{N}}$.
This is helpful since we now count the possibilities for $e_{2}\cdot\gamma$, a vector in the sublattice $N\mathbb{Z}\times\mathbb{Z}$. We do this by applying Lemma 2.3, which counts lattice points in balls. Since $z_{N}$ is balanced, the bound we obtain is roughly the volume of the ball $B(m^{1/2}\left\lVert e_{2}\right\rVert_{z})$ divided by the covolume of the lattice $z_{N}$. This gives $\ll m$ possibilities.
For $e_{1}\cdot\gamma$, we notice that $\left\lVert e_{1}\right\rVert_{z}$ is equal to $y+x/y\asymp 1/N+Nx$. Unfortunately, if $z$ is a balanced lattice, one can compute that we must have a bound $x\gg 1/\sqrt{N}$. Thus the norm above can be rather large. Even though $L_{z}$ is balanced, the size of the ball would give a hopelessly large bound.
@@
Since $y\asymp 1/N$, we see that if $m\ll N^{1-\varepsilon}$, the ball we obtain has a small radius of size $o(1)$. Since $L_{z}$ is a balanced lattice, we can only have at most one lattice point in such a small ball, regardless of its centre. For every vector $e_{2}\cdot\gamma$ fixed as above, this leaves at most one possibility for $e_{1}\cdot\gamma$. Therefore, the second row $e_{2}\cdot\gamma$ already fixes the whole matrix $\gamma$.
@@
To refine the process above, we only partially fix the second row of $\gamma$. This seems difficult to do in standard coordinates, that is, working with the exact entries of $\gamma$. Instead, we choose a reduced basis, $v_{1}$ and $v_{2}$, for the balanced lattice $L_{z_{N}}$. An upshot of Fricke reduction is that we can choose $v_{2}=e_{2}\cdot z$ (we already noticed above that $e_{2}\cdot z$ is a shortest vector in $L_{z_{N}}$).
@@
Applying the determinant to $\gamma_{1}=\gamma_{2}+c\cdot\operatorname{id}_{2}$ and assuming that $\gamma_{1}$ has a square determinant imply that $-c$ gives a solution to
@@ 6.2. The iterative strategy
Let $z=n(x)a(y)\in\operatorname{SL}_{n}(\mathbb{R})$ be a matrix in Iwasawa form. Let $\gamma\in\mathcal{M}_{n}(\mathbb{Z},N)$ with $\det\gamma=m$ and
@@
$z^{-1}\cdot\gamma\cdot z\cdot z^{T}\cdot\gamma^{T}\cdot z^{-T}=O(m^{2/n}).$ (6.2)
@@
$\left\lVert v_{i}\right\rVert_{z}\ll m^{1/n}\cdot d_{i},$ (6.3)
@@
The strategy for counting the number of matrices $\gamma$ is to iteratively count the number of possibilities for its rows. More precisely, we first count the number of possible $\gamma_{n}=v_{n}$ by a lattice point counting argument, that is Lemma 2.3, since $\gamma_{n}\in\mathbb{Z}^{n}$. For each such fixed possibility, we then count the number of possible $\gamma_{n-1}$ by using the condition on $v_{n-1}$ in (6.3). For this observe that
@@
where $\xi\in\mathbb{R}$ can be computed from the $x$-coordinates of $z$ (in fact, $\xi=x_{n-1,n}$). Thus, having fixed $\gamma_{n}$, the condition can be interpreted as saying that $\gamma_{n-1}$ is a lattice point inside a ball with shifted centre. We can use that the bounds in 2.3 are independent of the centre of the ball. In the results below, we ultimately choose $m$ small enough so that the ball can only contain one lattice point.
@@
To get the point $z$ into a good position for applying the strategy above, we make the reduction given by Proposition 5.2 and assume the second alternative in the statement. In this case, we study the properties of all lattices derived from $z$ relevant for this and the next sections. For the other alternative we use a bound derived from the Whittaker expansion, for which we refer to Section 7.1.
@@
Let $N$ be a prime and let $z\in\mathbb{H}$ reduce to a compactum $\Omega$. Assume that $z$ has Iwasawa coordinates
[…]
for $i=2,\ldots,n-1$, with $d=\alpha(z)=\left\lVert e_{n}\right\rVert_{z}$ and satisfies
$z\in\mathcal{L}(B,A)\cap\mathcal{L}^{\prime}(B,A).$
Then the successive minima of $z$ and $z^{\prime}$ are all $\asymp_{\Omega}1$ and the successive minima of $z_{N}$ and $(z^{\prime})_{N}$ are all $\asymp_{\Omega}N^{(n-1)/n}$.
@@
Throughout this proof all implied constants are allowed to depend on $\Omega$ and, implicitly, $n$. We call a lattice $L$ balanced if $\lambda_{1}\asymp d(L)^{1/n}$, where $\lambda_{1}\leq\ldots\leq\lambda_{n}$ are the successive minima of $L$. By Minkowski’s theorem (2.1), generalising Lemma 2.7, this is equivalent to $\lambda_{i}\asymp d(L)^{1/n}$ for all $i=1,\ldots,n$. This, together with Lemma 2.1 on the dual lattice and Lemma 2.2 on the successive minima of exterior products, implies that $L$ is balanced if and only if the dual $L^{\ast}$ is balanced. Note also that the property of being balanced is invariant under scaling. Computing the determinants, we thus aim to prove that $L_{z},L_{z^{\prime}},L_{z_{N}},L_{(z^{\prime})_{N}}$ are balanced lattices. That $L_{z}$ is balanced is part of the assumption (see again Lemma 2.7). We also compute from the Iwasawa coordinates and the fact that $\det(z)=1$ that
$d^{n}\asymp N^{n-1}=\det(z_{N}).$
Now since $z\in\mathcal{L}^{\prime}(B,A)$, Table 1 shows that the first successive minimum of $L_{z^{\prime}}^{\ast}$ is equal to $N^{-1+1/n}\alpha(z)\asymp 1$. Therefore $L_{z^{\prime}}^{\ast}$ is balanced and so is $L_{z^{\prime}}$. Next, compute explicitly that
$(z_{N})^{-T}=N^{-1}\operatorname{diag}(1,\ldots,1,N)z^{-T}=N^{-1+1/n}z^{\prime}.$
By the above, it follows that $L_{z_{N}}^{\ast}$ is balanced and so is $L_{z_{N}}$. We finally note that
$(z^{\prime})_{N}=N^{1-1/n}z^{-T}$
so the same reasoning implies that $L_{(z^{\prime})_{N}}$ is balanced. ∎
@@
Let $N$ be a prime and let $z\in\mathbb{H}$ reduce to a compactum $\Omega$. Assume that $z$ has Iwasawa coordinates
[…]
for $i=2,\ldots,n-1$, and $d=\alpha(z)=\left\lVert e_{n}\right\rVert_{z}$, and satisfies
$z\in\mathcal{L}(B,A)\cap\mathcal{L}^{\prime}(B,A).$
Then
|$|\{\gamma\in\mathcal{M}_{n}(\mathbb{Z},N)\mid\det(\gamma)\ll\Lambda^{n},\,z^{-1}\gamma z=O(\Lambda)\}|\ll_{n,\Omega}\Lambda^{n}(1+\Lambda^{n}/N)^{n-1}.$
@@
The bottom row $e_{n}\cdot z$ has congruence conditions and thus lies in the lattice corresponding to $z_{N}$. By Lemma 6.1, this is a balanced lattice, meaning that we can approximate all successive minima of $L_{z_{N}}$ by $\det(z_{N})^{1/n}=N^{(n-1)/n}\asymp d$. In fact, the proof of Lemma 6.1 shows that the minimum of $z_{N}$ is equal to the minimum of $N^{1-1/n}(z^{\prime})^{-T}$, which is $\alpha(z)=d$. Thus $e_{n}z_{N}$ is a vector of shortest length in $L_{z_{N}}$. Recall now the condition
@@
possibilities for the row $\gamma_{n}=e_{n}\cdot\gamma$. We continue bounding the number of possibilities for $\gamma_{i}$ inductively, $i<n$. More precisely, we suppose that $\gamma_{j}$ with $i<j\leq n$ are fixed. Then, by using the fact that $n(x)^{-1}$ is unipotent upper triangular in condition (6.3), the number of possibilities left for $\gamma_{i}$ is bounded by the number of lattice points in $L_{z}$ in a ball of radius $L\cdot d_{i}$ with fixed centre determined by the $\gamma_{j}$, $i<j$, and $n(x)$. Next, note that the successive minima of $L_{z}$ are all $\asymp_{n,\Omega}1$, by Lemma 2.7. Furthermore,
$d_{i}^{n}=(dy_{1}\cdots y_{n-i})^{n}\asymp_{n,\Omega}1/N.$
By Lemma 2.3, there are at most
$\ll_{n,\Omega}1+\Lambda d_{i}+\cdots+(\Lambda d_{i})^{n}\ll_{n,\Omega}1+\frac{\Lambda^{n}}{N}$
possibilities for $\gamma_{i}$. Putting all bounds together, we bound the number of matrices $\gamma$ by
$\ll_{n,\Omega}\Lambda^{n}(1+\Lambda^{n}/N)^{n-1}.$
@@
The last part of the proof above shows that, as long as $\Lambda$ is small enough in terms of $N$, the choice of last row of $\gamma$ already determines the whole matrix.555 We also remark that numerical experiments in dimension $n=2$ seem to indicate that the bound we obtain for the possibilities for the last row might be sharp.
@@ 6.3. Detecting determinants that are higher powers
Assume the same conditions as in Proposition 6.2. Additionally, let $L\ll N^{1/n^{2}-\varepsilon}$ and $N\gg_{\Omega}1$ be large enough. Then
@@
Let $\gamma\in H_{\ast}(z,m^{n},N)$, $m\asymp L^{n}$, and consider again the number of possibilities for the last row $\gamma_{n}$. For this, let $b_{1},\ldots,b_{n-1},e_{n}z$ be a reduced basis for $L_{z_{N}}$ (see Section 2.3 for the definition, which we apply to $N^{-(n-1)/n}z_{N}\in\operatorname{SL}_{n}(\mathbb{R})$). From the proof of Proposition 6.2, we note again that $e_{n}$ is a vector of shortest length in $L_{z_{N}}$, where
$\left\lVert e_{n}\right\rVert_{z}=\left\lVert e_{n}\right\rVert_{z_{N}}=d\asymp_{n,\Omega}N^{(n-1)/n}.$
By Minkowski’s theorem, we also have $\left\lVert b_{i}\right\rVert\asymp_{n,\Omega}N^{(n-1)/n}$. Now $\gamma_{n}\in L_{z_{N}}$, so it can be written as
$\gamma_{n}=\sum_{i=1}^{n-1}a_{i}b_{i}+a_{n}e_{n}z$
with $a_{i}\in\mathbb{Z}$. By Lemma 2.6 and recalling the condition $\left\lVert\gamma_{n}\right\rVert_{z}\ll L^{\nu}d$ from (6.3), we deduce that $a_{i}\ll L^{\nu}$, for all $1\leq i\leq n$. There are $L^{(n-1)\nu}$ possibilities for $a_{1},\ldots,a_{n-1}$. Choose any such combination of coefficients and assume there exist $\gamma\in H_{\ast}(z,m^{\nu},N)$ and $\gamma^{\prime}\in H_{\ast}(z,l^{\nu},N)$, for $m,l\asymp L^{n}$, such that $\gamma_{n}=\sum_{i=1}^{n-1}a_{i}b_{i}+a_{n}e_{n}z$ and
@@
Then $\lambda\in\mathbb{Z}$ and $\lambda\ll L^{\nu}$. Observe also that the matrix $\gamma-\gamma^{\prime}$ satisfies the same geometric conditions (6.3) as $\gamma$ and $\gamma^{\prime}$, simply by the triangle inequality (with a doubled implied constant, of course). We now apply the same iterative process as in the proof of Proposition 6.2. We note however that, under the present conditions, each step yields at most one possibility. Indeed, fix the last row of $\gamma-\gamma^{\prime}$, having the form $\lambda e_{n}$, by fixing $\lambda\ll L^{\nu}$. Next, the number of possibilities for the row $(\gamma-\gamma^{\prime})_{n-1}$ is bounded by the number of $L_{z}$-lattice points in a ball of radius $L^{\nu}\cdot N^{-1/n}$ centred at $x_{n-1,n}\cdot\lambda e_{n}z$, where $x_{n-1,n}$ is one of the $x$-coordinates of $z$. By assumption, the radius is bounded by $N^{-\varepsilon}$. However, if $N$ is large enough, this is greater than the first successive minimum of $z$, which is $\asymp_{\Omega}1$. There is thus only one possible lattice point. On the other hand, it is clear that the multiple $\lambda\cdot\operatorname{id}_{n}$ of the identity matrix lies in the set $H(z,\lambda,N)$. Since $\lambda\ll L^{\nu}$, we see that $\lambda e_{n-1}\cdot z$ satisfies the condition of the lattice point above (again, condition (6.3)). Consequently, it follows that
@@
Since this polynomial is defined over $\mathbb{Z}$ and irreducible over $\mathbb{Q}$ by assumption, we count the number of such solutions using Heath-Brown’s Theorem 3 in [HB02]. In the notation there, after homogenising the polynomial, we set $B_{1}=L^{\nu}$ for the bound on $\lambda$, then $B_{2}=L^{n}$ for the bound on $l$, and finally $B_{3}=1$ for the bound on the additional variable. Then we compute $T=L^{n\nu}$ and $V=L^{\nu+n}$. Heath-Brown’s result then gives the bound
@@
We are now left with counting degenerate matrices. This is reminiscent of treating the special case of parabolic matrices in [HT12, Lemma 2]. For this we restrict to prime degrees, allowing for a clean classification of the degenerate case.
Let $n\geq 2$ be prime. Since $\chi_{\gamma}$ is a polynomial of degree $n$ over $\mathbb{Q}$, a result of Schinzel [Sch67] shows that
@@
The method of proof in Proposition 6.4 provides even more. We recall at this point that the determinants $m^{\nu}$ appearing in the counting problem have a particular shape, namely $m=p\cdot q^{n-1}$, where $p$ and $q$ are primes of size $L$ (see the amplifier in Proposition 3.4). We are thus averaging over a set of size $L^{2}$. However, we can consider the special case $p=q$ to reduce this size.
@@
We follow the proof of Proposition 6.4, but first we fix the determinant $p^{n^{2}}$, where $p\asymp L$. There are, of course, at most $L$ such determinants. Now the number of choices for a potential last row of $\gamma\in H_{\text{par}}(z,p^{n^{2}},N)$ up to the contribution of $e_{n}$, i.e. up to the last component, is bounded by $L^{(n-1)n}$. Choose $\gamma$ and $\gamma^{\prime}$ two matrices in $H_{\text{par}}(z,p^{n^{2}},N)$ with the same last row up to the last component. As in the proof of Proposition 6.4, we find that

### v2

@@ 6.1. An overview
$z=\begin{pmatrix}\sqrt{y}&x/\sqrt{y}\\ 0&1/\sqrt{y}\end{pmatrix}=n(x)\cdot d\cdot\operatorname{diag}(y,1),$
where $d=1/\sqrt{y}$, and let $\gamma\in H(z,m,N)$. The bound
(6.1) $z^{-1}\gamma z=O(m^{1/2})$
@@
for $i=1,2$, where $B(r)$ is a Euclidean ball of radius $O(r)$ around 0.
We assume now that $z$ lies in bulk $\Omega_{N}$, i.e. it satisfies the conditions in case (2.b) of Proposition 5.2. In this context, we have that $L_{z}$ is a balanced lattice, and $1/N\ll y\ll 1/\sqrt{N}$. Importantly, the shortest length in $L_{z^{\prime}}$ is given by $d/\sqrt{N}$ (note either that rank 2 lattices are isometric to their duals or recall the discussion in Section 5.1). Let
@@
which defines a sublattice of index $N$ of $L_{z}$. In fact, using the classical definition of the Fricke involution and observing that
$\begin{pmatrix}0&-1\\ N&0\end{pmatrix}=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\cdot\begin{pmatrix}N&0\\ 0&1\end{pmatrix}$,
we easily see that $L_{z_{N}}$ is equal to $N^{1/2}L_{z^{\prime}}$. As such, the shortest length in $L_{z_{N}}$ is given by $d=\left\lVert e_{2}\right\rVert_{z}=\left\lVert e_{2}\right\rVert_{z_{N}}$.
This is helpful since we now count the possibilities for $e_{2}\cdot\gamma$, a vector in the sublattice $N\mathbb{Z}\times\mathbb{Z}$. We do this by applying Lemma 2.3, which counts lattice points in balls. Since the radius of the ball is given by $\sqrt{m}\cdot\left\lVert e_{2}\right\rVert_{z}$ and $\left\lVert e_{2}\right\rVert_{z}=d$ is the smallest length in $L_{z_{n}}$, the bound we obtain is
$1+\frac{d\sqrt{m}}{d}+\frac{d^{2}m}{d^{2}}\ll m,$
by bounding $\lambda_{1}\lambda_{2}\gg\lambda_{1}^{2}$.
For $e_{1}\cdot\gamma$, we notice that $\left\lVert e_{1}\right\rVert_{z}$ is equal to $y+x/y\asymp 1/N+Nx$. Unfortunately, if $z$ is a balanced lattice, one can compute that we must have a bound $x\gg 1/\sqrt{N}$. Thus, the norm above can be rather large. Even though $L_{z}$ is balanced, the size of the ball would give a hopelessly large bound.
@@
Since $y\ll 1/\sqrt{N}$, we see that if $m\ll N^{1/2-\varepsilon}$, the ball we obtain has a small radius of size $o(1)$. Since $L_{z}$ is a balanced lattice, we can only have at most one lattice point in such a small ball, regardless of its centre. For every vector $e_{2}\cdot\gamma$ fixed as above, this leaves at most one possibility for $e_{1}\cdot\gamma$. Therefore, the second row $e_{2}\cdot\gamma$ already fixes the whole matrix $\gamma$.
@@
To refine the process above, we only partially fix the second row of $\gamma$. This seems difficult to do in standard coordinates, that is, working with the exact entries of $\gamma$. Instead, we choose a reduced basis, $v_{1}$ and $v_{2}$, for the lattice $L_{z_{N}}$. An upshot of Fricke reduction is that we can choose $v_{2}=e_{2}\cdot z$ (we already noticed above that $e_{2}\cdot z$ is a shortest vector in $L_{z_{N}}$).
@@
Applying the determinant to $\gamma_{1}=\gamma_{2}+c\cdot\operatorname{id}_{2}$ and assuming that $\gamma_{1}$ has a square determinant imply that c-c gives a solution to
@@ 6.2. The iterative strategy
Let $z=n(x)a(y)\in\operatorname{SL}_{n}(\mathbb{R})$ be a matrix in Iwasawa form. We recall for the convenience of the reader that
$a(y)=\operatorname{diag}(d_{1},\ldots,d_{n})=d\cdot\begin{pmatrix}y_{1}\cdots y_{n-1}&&&\\ &\ddots&&\\ &&y_{1}&\\ &&&1\end{pmatrix}$.
Let $\gamma\in\mathcal{M}_{n}(\mathbb{Z},N)$ with $\det\gamma=m$ and
@@
(6.2) $z^{-1}\cdot\gamma\cdot z\cdot z^{T}\cdot\gamma^{T}\cdot z^{-T}=O(m^{2/n}).$
@@
(6.3) $\left\lVert v_{i}\right\rVert_{z}\ll m^{1/n}\cdot d_{i},$
@@
The strategy for counting the number of matrices $\gamma$ is to iteratively count the number of possibilities for its rows. More precisely, we first count the number of possible $\gamma_{n}=v_{n}$ by a lattice point counting argument using Lemma 2.3. For each such fixed possibility, we then count the number of possible $\gamma_{n-1}$ by using the condition on $v_{n-1}$ in (6.3). For this observe that
@@
where $\xi\in\mathbb{R}$ can be computed from the $x$-coordinates of $z$ (in fact, $\xi=x_{n-1,n}$). Thus, having fixed $\gamma_{n}$, the condition can be interpreted as saying that $\gamma_{n-1}$ is a lattice point inside a ball with shifted centre. We can use that the bounds in Lemma 2.3 are independent of the centre of the ball. In the results below, we ultimately choose $m$ small enough so that the ball can only contain one lattice point.
@@
To have $z$ in a good position for applying the strategy above, we make the reduction given by Proposition 5.2 and assume case (2.b) in the statement.
@@
Let $N$ be a prime and let $z\in\mathbb{H}$ satisfy the conditions of case (2.b) in Proposition 5.2. In particular, $d=\alpha(z)=\left\lVert e_{n}\right\rVert_{z}$ and $z\in\mathcal{L}^{\prime}(B,A)$. Then the $\lambda_{1}(L_{z_{n}})=d=\left\lVert e_{n}\right\rVert_{z_{N}}$.
@@
Since $z\in\mathcal{L}^{\prime}(B,A)$, we have $\lambda_{1}(L_{z^{\prime}}^{\ast})=dN^{-1+1/n}$. We now compute that
$(z^{\prime})^{-T}=A_{N}^{-1}z=N^{1/n}\cdot\operatorname{diag}(1,\ldots,1,N)^{-1}z=N^{-1+1/n}z_{N}.$
@@
Let $N$ be a prime and let $z\in\mathbb{H}$ satisfy the conditions of case (2.b) in Proposition 5.2. Then
$|\{\gamma\in\mathcal{M}_{n}(\mathbb{Z},N)\mid\det(\gamma)\ll\Lambda^{n},\,z^{-1}\gamma z=O(\Lambda)\}|\ll_{n,\Omega}\Lambda^{n}(1+\Lambda^{n}/\sqrt{N})^{n-1}.$
@@
The bottom row $e_{n}\cdot\gamma z$ has congruence conditions and thus lies in the lattice corresponding to $z_{N}$. By Lemma 6.1, the minimum of $L_{z_{N}}$ is equal $\alpha(z)=d$ and $e_{n}z_{N}$ is a vector of shortest length in $L_{z_{N}}$.
Recall now the condition
@@
possibilities for the row $\gamma_{n}=e_{n}\cdot\gamma$, using the trivial bounds $\lambda_{i}\geq\lambda_{1}$ for all $i$ for successive minima.
We continue bounding the number of possibilities for $\gamma_{i}$ inductively, $i<n$. More precisely, we suppose that $\gamma_{j}$ with $i<j\leq n$ are fixed. Then, by using the fact that $n(x)^{-1}$ is unipotent upper triangular in condition (6.3), the number of possibilities left for $\gamma_{i}$ is bounded by the number of lattice points in $L_{z}$ in a ball of radius $L\cdot d_{i}$ with fixed centre determined by the $\gamma_{j}$, $i<j$, and $n(x)$.
Next, note that $L_{z}$ is balanced, meaning that the successive minima of $L_{z}$ are all $\asymp_{n,\Omega}1$, as in Lemma 2.7. Furthermore,
$d_{i}^{n}=(dy_{1}\cdots y_{n-i})^{n}\ll_{n,\Omega}1/\sqrt{N},$
by assumption. By Lemma 2.3, there are at most
$\ll_{n,\Omega}1+\Lambda d_{i}+\cdots+(\Lambda d_{i})^{n}\ll_{n,\Omega}1+\frac{\Lambda^{n}}{\sqrt{N}}$
possibilities for $\gamma_{i}$.
Putting all bounds together, we bound the number of matrices $\gamma$ by
$\ll_{n,\Omega}\Lambda^{n}(1+\Lambda^{n}/\sqrt{N})^{n-1}.$
@@
The last part of the proof above shows that, as long as $\Lambda$ is small enough in terms of $N$, the choice of last row of $\gamma$ already determines the whole matrix. [^7] We also remark that numerical experiments in dimension $n=2$ seem to indicate that the bound we obtain for the possibilities for the last row might be sharp.
@@ 6.3. Detecting determinants that are higher powers
Let $N$ be a prime and let $z\in\mathbb{H}$ satisfy the conditions of case (2.b) in Proposition 5.2. Additionally, let $L\ll N^{1/2n^{2}-\varepsilon}$ and $N\gg_{\Omega}1$ be large enough. Then
@@
Let $\gamma\in H_{\ast}(z,m^{\nu},N)$, $m\asymp L^{n}$, and consider again the number of possibilities for the last row $\gamma_{n}$. For this, let $b_{1},\ldots,b_{n-1},e_{n}z$ be a reduced basis for $L_{z_{N}}$ (see Section 2.3 for the definition, which we apply to $N^{-(n-1)/n}z_{N}\in\operatorname{SL}_{n}(\mathbb{R})$). From Lemma 6.1, we note again that $e_{n}$ is a vector of shortest length in $L_{z_{N}}$, where $\left\lVert e_{n}\right\rVert_{z_{N}}=d$. We have $\left\lVert b_{i}\right\rVert\geq d$.
Now $\gamma_{n}z\in L_{z_{N}}$, so it can be written as
$\gamma_{n}z=\sum_{i=1}^{n-1}a_{i}b_{i}+a_{n}e_{n}z_{N}$
with $a_{i}\in\mathbb{Z}$. By Lemma 2.6 and recalling the condition $\left\lVert\gamma_{n}\right\rVert_{z}\ll L^{\nu}d$ from (6.3), we deduce that $a_{i}\ll L^{\nu}$, for all $1\leq i\leq n$.
There are $L^{(n-1)\nu}$ possibilities for $a_{1},\ldots,a_{n-1}$. Choose any such combination of coefficients and assume there exist $\gamma\in H_{\ast}(z,m^{\nu},N)$ and $\gamma^{\prime}\in H_{\ast}(z,l^{\nu},N)$, for $m,l\asymp L^{n}$, such that $\gamma_{n}=\sum_{i=1}^{n-1}a_{i}b_{i}+a_{n}e_{n}z$ and
@@
Then $\lambda\in\mathbb{Z}$ and $\lambda\ll L^{\nu}$. Observe also that the matrix $\gamma-\gamma^{\prime}$ satisfies the same geometric conditions (6.3) as $\gamma$ and $\gamma^{\prime}$, simply by the triangle inequality (with a doubled implied constant, of course).
We now apply the same iterative process as in the proof of Proposition 6.2. We note however that, under the present conditions, each step yields at most one possibility. Indeed, fix the last row of $\gamma-\gamma^{\prime}$, having the form $\lambda e_{n}$, by fixing $\lambda\ll L^{\nu}$. Next, the number of possibilities for the row $(\gamma-\gamma^{\prime})_{n-1}$ is bounded by the number of $L_{z}$-lattice points in a ball of radius $L^{\nu}\cdot N^{-1/2n}$ centred at $x_{n-1,n}\cdot\lambda e_{n}z$, where $x_{n-1,n}$ is one of the $x$-coordinates of $z$. By assumption, the radius is bounded by $N^{-\varepsilon}$. However, if $N$ is large enough, this is greater than the first successive minimum of $z$, which is $\asymp_{\Omega}1$. There is thus only one possible lattice point.
On the other hand, it is clear that the multiple $\lambda\cdot\operatorname{id}_{n}$ of the identity matrix lies in the set $H(z,\lambda^{n},N)$. Since $\lambda\ll L^{\nu}$, we see that $\lambda e_{n-1}\cdot z$ satisfies the condition of the lattice point above (again, condition (6.3)). Consequently, it follows that
@@
Since this polynomial is defined over $\mathbb{Z}$ and irreducible over $\mathbb{Q}$ by assumption, we count the number of such solutions using Heath-Brown’s Theorem 3 in [23]. In the notation there, after homogenising the polynomial, we set $B_{1}=L^{\nu}$ for the bound on $\lambda$, then $B_{2}=L^{n}$ for the bound on $l$, and finally $B_{3}=1$ for the bound on the additional variable. Then we compute $T=L^{n\nu}$ and $V=L^{\nu+n}$. Heath-Brown’s result then gives the bound
@@
We are now left with counting degenerate matrices. This is reminiscent of treating the special case of parabolic matrices in [21, Lemma 2]. For this we restrict to prime degrees, allowing for a clean classification of the degenerate case.
Let $n\geq 2$ be prime. Since $\chi_{\gamma}$ is a polynomial of degree $n$ over $\mathbb{Q}$, a result of Schinzel [37] shows that
@@
The method of proof in Proposition 6.4 provides even more. We recall at this point that the determinants $m^{\nu}$ appearing in the counting problem have a particular shape, namely $m=p\cdot q^{n-1}$, where $p$ and $q$ are primes of size $L$ (see the amplifier in Proposition 3.4). We are thus averaging over a set of size $L^{2}$. However, we can consider the special case $p=q$, which gives a smaller number of terms.
@@
We follow the proof of Proposition 6.4, but first we fix the determinant $p^{n^{2}}$, where $p\asymp L$. There are, of course, at most $L$ such determinants. Now the number of choices for a potential last row of $\gamma\in H_{\text{par}}(z,p^{n^{2}},N)$ up to the contribution of $e_{n}$, i.e. up to the last component, is bounded by $L^{(n-1)n}$. Choose $\gamma$ and $\gamma^{\prime}$ two matrices in $H_{\text{par}}(z,p^{n^{2}},N)$ with the same last row up to the last component.
As in the proof of Proposition 6.4, we find that